OpenStudy (anonymous):

how to simplify ((16x^4)^1/4)?

OpenStudy (anonymous):

\[(16x^4)^.25\]

OpenStudy (aum):

\[\large (16x^4)^{1/4} = (2^4*x^4)^{1/4} = ((2x)^4)^{1/4} = (2x)^{4*1/4} = 2x^1 = 2x \]

OpenStudy (anonymous):

\[(2^{4/4}x^{4/4})\]

OpenStudy (anonymous):

Thanks to all of you.You been a great help :)